OpenStudy (anonymous):

What is the equation in slope-intercept form of the line through the points (2, -3) and (-3, 7)?

OpenStudy (anonymous):

first find slope m= 7+3/-3-2=10/-5=-2

OpenStudy (anonymous):

slope intercept form y=mx + b

OpenStudy (anonymous):

I'm confused. What's the answer?

OpenStudy (anonymous):

Well first you should determine the slope. Here's how to remember: "Slope is easy and it's fun. Just remember rise over run!" From (-3, 7) to (2, -3), the line goes down 10 and over 5, so the slope is -10/5 = -2. So your M is -2. Now just pick a point, substitute the x and y into your y=mx+b equation and solve for b. So picking (2, -3) gives: y= m*x + b -3 = -2*2+b 1 = b So your equation is y = -2*x + 1